OpenStudy (anonymous):

Find the x and y values: x-2y=3, x-y=-7

OpenStudy (tkhunny):

Have you considered subtracting the second equation from the first?

OpenStudy (anonymous):

How would I do that?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{x-2y=3---A\\ x-y=-7----B }\end{align}\) multiply \(\Large B\) by \(\large -1\) \(\large\tt \begin{align} \color{black}{ -(x-y)=-(-7)----B\\ -x+y=7--C}\end{align}\) now add\(\large A\) and \(\large C\) \(\large\tt \begin{align} \color{black}{x-2y=3---A\\ -x+y=7--C\\ (x-2y)+(-x+y)=(3)+(7)\\ x-2y-x+y=3+7\\ -y=10\\ \huge y=-10}\end{align}\) now substitute value of \(\large y=-10\) in \(\Large C \) or \(\large B\) and get the value of \(\large x\)

OpenStudy (tkhunny):

You should be up on subtraction. Just combine the like terms.

OpenStudy (anonymous):

10-7=3

OpenStudy (tkhunny):

x-x = 0 -2y + y = -y 3 + 7 = 10 So, -y = 10 and y = -10 Substitute this value into either original equation and find x.